OpenStudy (anonymous):

Solve by completing the square. x2+6x−16=0 x=__ or x=__

OpenStudy (calculusxy):

\(x^2 + 6x - 16 = 0\) \(x^2 + 6x - 16 + 16 = 0 + 16\) \(x^2 + 6x + (\frac{6}{2})^2 = 16 + (\frac{6}{2})^2\) \(x^2 + 6x + 3^2 = 16 + 3^2\) \(x^2 + 6x + 9 = 16 + 9\) \(\sqrt{x^2 + 6x + 9} = \sqrt{25}\) \((x + 3)^2 = -5 \text{ } \text{or} \text{ } 5\) solve for x in these two cases \(x+ 3 = 5\) and \(x + 3 = -5\)

OpenStudy (welshfella):

X^2+6x-16=0 (X+3)^2-9-16=0 (X+3)^2 = 25 Can u continue.?

OpenStudy (anonymous):

mates got it

OpenStudy (calculusxy):

what are the two values (just to confirm) ?

OpenStudy (anonymous):

you nailed it

OpenStudy (anonymous):

good show

OpenStudy (calculusxy):

what is this supposed to be about, may i know ?

OpenStudy (anonymous):

solving by completing the squar

OpenStudy (anonymous):

thats not right one moment

OpenStudy (anonymous):

cumpass says it is squar in a quadratic

OpenStudy (anonymous):

square

OpenStudy (calculusxy):

you mean to use the quadratic formula?

OpenStudy (anonymous):

yes thank you

OpenStudy (calculusxy):

it's the same concept except you need to know where the a, b, and c values are in a quadratic expression. do you know where they are? hint: ax^2 + bx + c

OpenStudy (calculusxy):

and after that just plug those values into the formula: \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

oh